Circular No. 5645

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1992bg AND 1992bh IN ANONYMOUS GALAXIES

J. Maza, University of Chile; and M. Hamuy, Cerro Tololo Interamerican Observatory, report the discovery by R. Antezana, University of Chile, of two supernovae on 20-min unfiltered IIa-O plates taken by G. Valladares with the CTIO Curtis Schmidt telescope. SN 1992bg (m_{pg} ∼ 17.8 on the Oct. 21.34 UT discovery plate) is located ∼ 3′′ west and 6′′ north of the nucleus of a spiral galaxy at α = 7^{h}41^{m}19^{s}.0, δ = −62^{o}24′06′′ (equinox 1950.0). SN 1992bh (m_{pg} ∼ 17.3 on the Oct. 21.32 discovery plate) is located ∼ 1′′ east and 2′′.5 south of the nucleus of a spiral galaxy at α = 4^{h}58^{m}40^{s}.0, δ = −58^{o}54′10′′. Confirmation of both supernovae was made by M. Navarrete (CTIO) on CCD B, V, and I images obtained on Oct. 24.31 with the CTIO 4-m telescope. Spectrograms (range 320-750 nm) of both supernovae were obtained by Hamuy on Oct. 27.2 with the CTIO 1.5-m telescope, showing the characteristic Si II 635.5-nm feature of type-Ia supernovae. For SN 1992bg, the minimum of the Si II absorption yields an expansion velocity of ∼ 11 000 km/s with respect to the galaxy’s rest frame; this spectrum shows resemblance to that of the type-Ia SN 1992A taken 5 days past maximum light. The spectrum of the nucleus of the parent galaxy seems to yield a redshift of z = 0.035. For SN 1992bh, the minimum of the Si II absorption yields an expansion velocity of ∼ 12 000 km/s with respect to the galaxy’s rest frame; this spectrum shows resemblance to that of the type-Ia SN 1992A taken at maximum light. The redshift inferred from Hα emission of the underlying galaxy is z = 0.044.

PERIODIC COMET SWIFT-TUTTLE (1992t) S. Nakano, Sumoto, Japan, reports that Y. Kushida, Yatsugatake

South Base Obesrvatory, obtained photographs with a 0.25-m f/3.4 reflector, showing a faint tail ∼ 2^{o} and 1^{o}.6 long on Oct. 25.41 and 26.39 UT, respectively, toward the north-northeast. There was no such long tail on Oct. 21, but there was an indication of tail growing on Oct. 22.41.

Total visual magnitude estimates (cf. IAUC 5636): Oct. 17.99 UT, 6.9 (J. E. Bortle, Stormville, NY, 10×50 binoculars); 19.79, 6.9 (S. Canonaco, Genk, Belgium, 10×50 binoculars); 22.99, 6.7 (G. R. Chester, Rixeyville, VA, 9×63 binoculars); 23.86, 6.4 (J. D. Shanklin, Cambridge, England, 10×80 binoculars); 25.11, 6.3 (C. S. Morris, near Pt. Conception, CA, 10×50 binoculars); 26.74, 6.5 (S. Baroni, Milan, Italy, 20×80 binoculars); 28.01, 6.1 (D. W. E. Green, Harvard, MA, 12×50 binoculars).

1992 October 28 Daniel W. E. Green